OpenStudy (anonymous):

2y = -3x + 8 is an equation that represents a line parallel to the line 6x + 2y = 9. (True or False)

OpenStudy (jhannybean):

solve for y in the second equation by subtracting -6x from both sides \[6x+2y = 9 \to -6x+6x + 2y =9-6x\]\[2y= -6x+9\] solve for y \[y=\frac{-6}{2}x+\frac{9}{2}\]

OpenStudy (jhannybean):

then solve for y in the other equation \[2y = -3x+8 \to y=\frac{-3}{2}x+ \frac{8}{2}\]

OpenStudy (anonymous):

so it will be false? right?

OpenStudy (jhannybean):

you have slopes -3 and -3/2 In parallel lines the slopes are the same, are these slopes the same to you?

OpenStudy (jhannybean):

You are right :D

OpenStudy (anonymous):

slope of parellel lines must be same first line's slope is -3/2 seconde line's slope is -6/2 they aren't same so they are not parellel

OpenStudy (jhannybean):

Good job.

OpenStudy (anonymous):

THANK YALL SO MUCH!! HAHA

OpenStudy (jhannybean):

:) no problem.